# Anisotropic Quantum Annealing vs Trit Annealing

### v1

@@ Abstract
Quantum annealing offers a promising strategy for solving complex optimization problems by encoding the solution into the ground state of a problem Hamiltonian. While most implementations rely on spin-$1/2$ systems, we explore the performance of quantum annealing on a spin-$1$ system where the problem Hamiltonian includes a single ion anisotropy term of the form $D\sum(S^{z})^{2}$. Our results reveal that for a suitable range of the anisotropy strength $D$, the spin-$1$ annealer reaches the ground state with higher fidelity. We attribute this performance to the presence of the intermediate spin level and the tunable anisotropy, which together enable the algorithm to traverse the energy landscape through smaller, incremental steps instead of a single large spin flip. This mechanism effectively lowers barriers in the configuration space and stabilizes the evolution. These findings suggest that higher spin annealers offer intrinsic advantages for robust and flexible quantum optimization, especially for problems naturally formulated with ternary decision variables.
@@ I Introduction
Combinatorial optimization arises across physics, computer science, and engineering, with applications from materials and circuit design to logistics, portfolio optimization, and scheduling Harrison2022max3cutcoherent; Weggemans2022solvingcorrelation. Many such problems can be cast as minimizing a cost function over discrete variables, naturally mapping to the ground state of an effective Ising-type Hamiltonian with local fields and pairwise couplings Finnila1994; Kadowaki1998; Farhi2001; Das2008; Lucas2014. Candidate solutions correspond to spin configurations, while optimal or near-optimal solutions lie at low energies. Since the configuration space grows exponentially with system size, classical methods can struggle on rugged, frustrated landscapes with many nearly degenerate local minima.
@@
Quantum annealing (QA) is a heuristic approach to such optimization problems that leverages quantum fluctuations Kadowaki1998; Farhi2001; Albash2018Review. The protocol initializes the system in the ground state of a simple driver Hamiltonian—typically a transverse field term that does not commute with the problem Hamiltonian—and then interpolates toward the problem Hamiltonian under a time-dependent schedule. In the coherent, adiabatic limit, the system follows the instantaneous ground state provided the evolution is sufficiently slow. In practice, the runtime is constrained by the minimum spectral gap encountered along the annealing path Farhi2001; Albash2018Review; Jansen2006BoundsFT; small gaps therefore require long schedules and increase susceptibility to nonadiabatic excitations, decoherence, and control errors Bottarelli2025symmetrycd.
@@
Most theoretical and experimental QA platforms use effective spin-$1/2$ degrees of freedom (qubits), which directly encode binary decision variables Das2008; Albash2018Review; Hauke2020; Johnson2011; DenchevPRX2016; BoixoNC2016. However, many practically relevant optimization problems are intrinsically multivalued, motivating approaches that go beyond binary encoding. Examples include ternary decisions (e.g., buy/hold/sell) and formulations with variables taking values in sets such as $\{-1,0,+1\}$. On qubit based architectures, representing multivalued logic typically requires embedding each variable into multiple qubits together with penalty terms enforcing logical constraints Weggemans2022solvingcorrelation. These embeddings increase hardware overhead, reduce effective connectivity, and complicate the energy landscape, potentially shrinking relevant gaps and amplifying sensitivity to noise and control imperfections. Higher spin systems provide a natural alternative in which multilevel quantum degrees of freedom directly encode multivalued variables.
@@
Spin-11 systems, in particular, have three eigenstates of $S^{z}$ labeled by $m_{s}=-1,0,+1$, enabling a direct representation of ternary variables without elaborate embeddings. This has motivated interest in multilevel quantum optimization and in understanding how extra levels affect adiabatic protocols Amin2013Qudits; WangHuSandersKais2020; Weggemans2022solvingcorrelation; Bottarelli2025symmetrycd. Beyond compact encoding, spin-$1$ models also introduce control parameters absent in spin-$1/2$ systems. A key example is the single ion anisotropy term $D\sum_{i}(S_{i}^{z})^{2}$, familiar from spin-$1$ magnetism such as the Blume–Capel model Blume1966; Capel1966, which tunes the relative energetics of the $m_{s}=0$ and $m_{s}=\pm 1$ states. Since $(S^{z})^{2}$ is proportional to the identity for spin-$1/2$, this anisotropy provides a genuine additional control knob for spin-$1$ annealers and gives rise to richer behavior along the annealing path.
@@ II Theoretical Framework
  • $D$ is the single ion anisotropy parameter,
@@
with gap $\Delta(t)=E_{1}(t)-E_{0}(t)$. Near an avoided crossing with minimum gap $\Delta_{\min}$, a two level linearization yields the Landau–Zener formula Landau
@@
The corresponding classical configuration space is
@@
At temperature $T>0$, a single-site Metropolis update Metropolis1953 proceeds as follows:
@@ III Methods
TA is defined on the same set of spin configurations by taking the classical cost function to be the diagonal of the quantum problem Hamiltonian, so that classical and quantum instances are exactly aligned at the level of their energy landscapes. One sweep comprises $N$ single site updates, in which each spin is updated once on average according to the local Boltzmann weights at the current effective temperature. The inverse temperature is reduced following the same schedule shape $g(t)$ used in the quantum protocol, now applied over the course of $S$ sweeps. For each disorder instance (coupling realization) we perform $R$ independent TA runs starting from random initial spin configurations drawn uniformly from the $3^{N}$ classical states. At the end of each run $(r)$ we record the terminal energy $E_{\mathrm{final}}^{(r)}$. The exact classical ground state energy $E_{0}$ is obtained by exhaustive enumeration of all configurations. We define the TA success probability as
@@
where $\mathbb{I}[\cdot]$ denotes the indicator function. The small numerical tolerance accounts for floating point round off while effectively requiring that the algorithm reaches the true ground state. By matching instance sets, schedule shapes, and problem Hamiltonians between AQA and TA, we can directly compare their success probabilities.
@@ IV Results
Figure 1: AQA versus TA across the coupling–anisotropy plane. Each panel displays the difference $P_{\mathrm{AQA}}-P_{\mathrm{TA}}$ over the $(J,D)$ plane for a fixed transverse field scale $e$ and annealing schedule $g(t)$, at longitudinal field $h=0.2$. Columns (left to right) correspond to schedules with late-time decay $g(t)\propto 1/\log(1{+}t)$, $1/\sqrt{t}$, $1/t$, and $1/t^{2}$, while rows (top to bottom) correspond to $e=1,5,10,20$. Red (blue) indicates parameter regions where AQA yields higher (lower) success probability than TA, and yellow markers indicate cases with $P_{\mathrm{AQA}}>0.9$. The dashed gray lines at $J=0$ and $D=0$ delineate the easy-plane ($D>0$) and easy-axis ($D<0$) sectors and assist in identifying symmetry-related structures.
Figure 1 compares the performance of AQA and TA across the $(J,D)$ plane for all transverse field scales and schedules considered. The dominant trends are governed by the driver strength $e$, the anisotropy sector (easy-plane $D>0$ versus easy-axis $D<0$), and the late-time decay of the schedule $g(t)$.
For weak driving ($e=1$, top row), TA is comparable to AQA over nearly the entire parameter range, as indicated by the predominance of blue regions across all schedules. In this regime, the transverse field contribution is insufficient, within the fixed runtime, to reliably generate transitions that improve the final success probability. The sparse red features occur in narrow parameter bands near boundaries between distinct classical ordering tendencies, consistent with cases where near degeneracies permit limited benefit from coherent mixing.
As the driver strength increases ($e=5$ and above), extended parameter regions emerge in which AQA outperforms TA. This advantage is most apparent in the easy-plane sector ($D>0$) at intermediate coupling magnitudes $|J|$, where red domains become substantial for $e=5$ and expand further for $e=10$ and $e=20$. The growth of these AQA favorable regions indicates that sufficiently strong transverse driving can more effectively exploit low energy structure associated with competing classical patterns.
@@
At fixed e, performance exhibits a systematic dependence on the schedule. Slower late-time decays expand the parameter regions where AQA outperforms TA, with the effect most pronounced at larger e. In contrast, faster decays yield comparatively larger regions where TA remains dominant. These observations are consistent with the view that, at fixed total runtime, emphasizing late-time evolution enhances AQA’s ability to resolve and traverse competing low-energy structures.
@@ V Conclusion
This study investigated quantum and classical annealing dynamics for an open boundary spin-$1$ Ising chain with quadratic anisotropy. Across the $(J,D)$ plane, we observed a consistent pattern: quantum advantage emerges when two conditions are jointly satisfied, namely $(i)$ a sufficiently delocalized initial state (large $e$) and $(ii)$ schedules that allocate more time at small $g$. In these regimes, especially near order-change ridges where the dominant pattern shifts (all-0, all-$\pm 1$, antiferromagnetic), the dynamics benefit from stepwise rearrangements rather than a single collective switch. This distributes the evolution over multiple avoided crossings and can improve adiabatic transport relative to schedules that effectively force a single collective switch.
@@
A second, schedule dependent effect concerns how the annealing time is distributed between the driver dominated and problem dominated parts of the evolution. For decays that naturally dwell near the final Hamiltonian (e.g., $g(t)\propto 1/\log(1{+}t)$) and at larger $e$, we observe systematically larger regions in which AQA outperforms TA, indicating that additional late-time budget is crucial for resolving competition between low energy patterns once the transverse field is weak. Linear ramps, by contrast, retain sizable TA dominant patches even at e=20e{=}20, underscoring that the terminal slowdown is an operational rather than purely cosmetic feature. Overall, TA performance follows the structure of classical barriers, while AQA shows additional sensitivity to schedule shape and transverse scale, consistent with coherent pathways that circumvent the need for collective flips (see App. A for supporting classical landscape diagnostics).
@@
All comparisons use matched time budgets and identical $g(t)$ grids, so the observed advantage is a finite time effect under controlled conditions rather than an asymptotic separation. The present study nevertheless considers a single chain with open boundary condition, a fixed symmetry breaking field $h$, and one classical baseline. Future comparisons against additional classical heuristics (e.g., simulated annealing with optimized temperature schedules, population annealing, or cluster updates adapted to ternary variables) would strengthen the baseline. Likewise, scaling with system size and alternative graph topologies (rings, weak disorder, or sparse long-range couplings) will be important to assess robustness. Finally, schedule design itself is a practical lever: local adiabatic ramps, explicit end pauses, or counter-diabatic corrections could sharpen the quantum protocol without altering the problem Hamiltonian.
@@
To the best of our knowledge, this is one of the first detailed investigations of ternary optimization within a quantum annealing framework using a spin-$1$ system with a quadratic anisotropy. We find evidence for a narrow but reproducible region in the $(J,D)$ plane where quantum systems outperform classical systems under matched finite time budgets. The advantage becomes most visible when the initial transverse scale $e$ is large and the schedule allocates sufficient time at small $g$, conditions that jointly promote coherent transport across sequences of avoided crossings near order-change boundaries, particularly in the easy-plane sector ($D>0$). Conversely, with insufficient driving or an aggressive terminal ramp, AQA can underperform even on classically easy instances.

### v2

@@ Abstract
Quantum annealing offers a promising strategy for solving complex optimization problems by encoding the solution into the ground state of a problem Hamiltonian. While most implementations rely on spin-$1/2$ systems, we explore the performance of quantum annealing on a spin-$1$ system where the problem Hamiltonian includes a single-ion anisotropy term of the form $D\sum(S^{z})^{2}$ and compare it with trit based simulated annealing. Our results reveal that for a suitable range of the anisotropy strength $D$, the spin-$1$ annealer reaches the ground state with higher success probability. We attribute this performance to the presence of the intermediate spin level and the tunable anisotropy, which together enable the algorithm to traverse the energy landscape through smaller, incremental steps instead of a single large spin flip. This mechanism effectively lowers barriers in the configuration space and stabilizes the evolution. These findings suggest that higher-spin annealers can provide a finite-time performance benefit relative to the trit annealing baseline studied here, indicating a potentially robust and flexible approach to quantum optimization for problems naturally formulated with ternary decision variables.
@@ I Introduction
Combinatorial optimization arises across physics, computer science, and engineering, with applications from materials and circuit design to logistics, portfolio optimization, and scheduling Harrison et al. (2022); Weggemans et al. (2022). Many such problems can be cast as minimizing a cost function over discrete variables, naturally mapping to the ground state of an effective Ising-type Hamiltonian with local fields and pairwise couplings Finnila et al. (1994); Kadowaki and Nishimori (1998); Childs et al. (2001); Das and Chakrabarti (2008); Lucas (2014). Candidate solutions correspond to spin configurations, while optimal or near-optimal solutions lie at low energies. Since the configuration space grows exponentially with system size, classical methods can struggle on rugged, frustrated landscapes with many nearly degenerate local minima.
@@
Quantum annealing (QA) is a heuristic approach to such optimization problems that leverages quantum fluctuations Kadowaki and Nishimori (1998); Childs et al. (2001); Albash and Lidar (2018). The protocol initializes the system in the ground state of a simple driver Hamiltonian—typically a transverse field term that does not commute with the problem Hamiltonian—and then interpolates toward the problem Hamiltonian under a time-dependent schedule. In the coherent, adiabatic limit, the system follows the instantaneous ground state provided the evolution is sufficiently slow. In practice, the runtime is constrained by the minimum spectral gap encountered along the annealing path Childs et al. (2001); Albash and Lidar (2018); Jansen et al. (2007); small gaps therefore require long schedules and increase susceptibility to nonadiabatic excitations, decoherence, and control errors Bottarelli et al. (2025).
@@
Most theoretical and experimental QA platforms use effective spin-$1/2$ degrees of freedom (qubits), which directly encode binary decision variables Das and Chakrabarti (2008); Albash and Lidar (2018); Hauke et al. (2020); Johnson et al. (2011); Denchev et al. (2016); Boixo et al. (2016). However, many practically relevant optimization problems are intrinsically multivalued, motivating approaches that go beyond binary encoding. Examples include ternary decisions (e.g., buy/hold/sell) and formulations with variables taking values in sets such as $\{-1,0,+1\}$. On qubit based architectures, representing multivalued logic typically requires embedding each variable into multiple qubits together with penalty terms enforcing logical constraints Weggemans et al. (2022). These embeddings increase hardware overhead, reduce effective connectivity, and complicate the energy landscape, potentially shrinking relevant gaps and amplifying sensitivity to noise and control imperfections. Higher spin systems provide a natural alternative in which multilevel quantum degrees of freedom directly encode multivalued variables.
@@
Spin-11 systems, in particular, have three eigenstates of $S^{z}$ labeled by $m_{s}=-1,0,+1$, enabling a direct representation of ternary variables without elaborate embeddings. This has motivated interest in multilevel quantum optimization and in understanding how extra levels affect adiabatic protocols Amin et al. (2013); Wang et al. (2020); Weggemans et al. (2022); Bottarelli et al. (2025). Beyond compact encoding, spin-$1$ models also introduce control parameters absent in spin-$1/2$ systems. A key example is the single-ion anisotropy term $D\sum_{i}(S_{i}^{z})^{2}$, familiar from spin-$1$ magnetism such as the Blume–Capel model Blume (1966); Capel (1966), which tunes the relative energetics of the $m_{s}=0$ and $m_{s}=\pm 1$ states. Since $(S^{z})^{2}$ is proportional to the identity for spin-$1/2$, this anisotropy provides a genuine additional control knob for spin-$1$ annealers and gives rise to richer behavior along the annealing path.
@@ II Theoretical Framework
  • $D$ is the single-ion anisotropy parameter,
@@
with gap $\Delta(t)=E_{1}(t)-E_{0}(t)$. Near an avoided crossing with minimum gap $\Delta_{\min}$, a two level linearization yields the Landau–Zener formula Zener (1932)
@@
We use the term trit for a classical three-valued variable $s\in\{-1,0,+1\}$. The corresponding classical configuration space is
@@
At temperature $T>0$, a single-site Metropolis update Metropolis et al. (1953) proceeds as follows:
@@ III Methods
TA is defined on the same set of spin configurations by taking the classical cost function to be the diagonal of the quantum problem Hamiltonian, so that classical and quantum instances are exactly aligned at the level of their energy landscapes. One sweep comprises $N$ single-site updates, in which each spin is updated once on average according to the local Boltzmann weights at the current effective temperature. The inverse temperature is reduced following the same schedule shape $g(t)$ used in the quantum protocol, now applied over the course of $S$ sweeps. For each instance, we perform $R$ independent TA runs starting from random initial spin configurations drawn uniformly from the $3^{N}$ classical states. At the end of each run $(r)$ we record the terminal energy $E_{\mathrm{final}}^{(r)}$. The exact classical ground state energy $E_{0}$ is obtained by exhaustive enumeration of all configurations. We define the TA success probability as
@@
where $\mathbb{I}[\cdot]$ denotes the indicator function. The small numerical tolerance accounts for floating-point round-off while effectively requiring that the algorithm reaches the true ground state.
@@ IV Results
Figure 1: AQA versus TA across the coupling–anisotropy plane. Each panel displays the difference $P_{\mathrm{AQA}}-P_{\mathrm{TA}}$ over the $(J,D)$ plane for a fixed transverse field scale $c$ and annealing schedule $g(t)$, at longitudinal field $h=0.2$. Columns (left to right) correspond to schedules with late-time decay $g(t)\propto 1/\log(1{+}t)$, $1/\sqrt{t}$, $1/t$, and $1/t^{2}$, while rows (top to bottom) correspond to $c=1,5,10,20$. Red (blue) indicates parameter regions where AQA yields higher (lower) success probability than TA, and yellow markers indicate cases with $P_{\mathrm{AQA}}>0.9$. The $(J,D)$ plane is sampled on a uniform grid with spacing $\Delta J=\Delta D=0.2$ over $J,D\in[-5,5]$. For each grid point, AQA is run for total runtime $T=1000$, and TA is run with an aligned budget of $S=1000$ sweeps and $R=20$ independent restarts. The dashed gray lines at $J=0$ and $D=0$ separate the easy-plane ($D>0$) and easy-axis ($D<0$) sectors and assist in identifying symmetry-related structures.
Figure 1 compares the performance of AQA and TA across the $(J,D)$ plane for all transverse field scales and schedules considered. Yellow markers highlight instances with $P_{\mathrm{AQA}}>0.9$; this threshold is used only to visually identify near-unit success and does not affect the underlying heatmaps. The dominant trends are governed by the driver strength $c$, the anisotropy sector (easy-plane $D>0$ versus easy-axis $D<0$), and the late-time decay of the schedule $g(t)$. We stress that throughout the paper the statement “AQA outperforms TA” is used in a strictly operational sense, i.e., $P_{\mathrm{AQA}}>P_{\mathrm{TA}}$ for the particular TA baseline considered here (local single-spin Metropolis updates for trits) under matched finite-time budgets, and should not be interpreted as a general separation from all classical annealing heuristics.
For weak driving ($c=1$, top row), TA is comparable to AQA over nearly the entire parameter range, as indicated by the predominance of blue regions across all schedules. In this regime, the transverse field contribution is insufficient, within the fixed runtime, to reliably generate transitions that improve the final success probability. The sparse red features occur in narrow parameter bands near boundaries between distinct classical ordering tendencies, consistent with cases where near degeneracies permit limited benefit from coherent mixing.
As the driver strength increases ($c=5$ and above), extended parameter regions emerge in which AQA outperforms TA. This advantage is most apparent in the easy-plane sector ($D>0$) at intermediate coupling magnitudes $|J|$, where red domains become substantial for $c=5$ and expand further for $c=10$ and $c=20$. The growth of these AQA favorable regions indicates that sufficiently strong transverse driving can more effectively exploit low energy structure associated with competing classical patterns.
@@
At fixed $c$, performance exhibits a systematic dependence on the schedule. Slower late-time decays expand the parameter regions where AQA outperforms TA, with the effect most pronounced at larger $c$. In contrast, faster decays yield comparatively larger regions where TA remains dominant. These observations are consistent with the view that, at fixed total runtime, emphasizing late-time evolution enhances AQA’s ability to resolve and traverse competing low-energy structures.
@@ V Conclusion
This study investigated quantum and classical annealing dynamics for an open boundary spin-$1$ Ising chain with quadratic anisotropy. Across the $(J,D)$ plane, we observed a consistent pattern: quantum advantage emerges when two conditions are jointly satisfied, namely $(i)$ a sufficiently delocalized initial state (large $c$) and $(ii)$ schedules that allocate more time at small $g$. In these regimes, especially near order-change ridges where the dominant pattern shifts (all-0, all-$\pm 1$, antiferromagnetic), the dynamics benefit from stepwise rearrangements rather than a single collective switch. This distributes the evolution over multiple avoided crossings and can improve adiabatic transport relative to schedules that effectively force a single collective switch.
@@
A second, schedule dependent effect concerns how the annealing time is distributed between the driver dominated and problem dominated parts of the evolution. For decays that naturally dwell near the final Hamiltonian (e.g., $g(t)\propto 1/\log(1{+}t)$) and at larger $c$, we observe systematically larger regions in which AQA outperforms TA, indicating that additional late-time budget is crucial for resolving competition between low energy patterns once the transverse field is weak. Linear ramps, by contrast, retain sizable TA dominant patches even at c=20c{=}20, underscoring that the terminal slowdown is an operational rather than purely cosmetic feature. Overall, TA performance follows the structure of classical barriers, while AQA shows additional sensitivity to schedule shape and transverse scale, consistent with coherent pathways that circumvent the need for collective flips (see App. A for supporting classical landscape diagnostics).
@@
All comparisons use matched time budgets and identical $g(t)$ grids, so the observed advantage is a finite-time effect under controlled conditions rather than an asymptotic separation. The present study nevertheless considers a single chain with open boundary condition, a fixed symmetry breaking field $h$, and one classical baseline. Future comparisons against additional classical heuristics (e.g., simulated annealing with optimized temperature schedules, population annealing, or cluster updates adapted to ternary variables) would strengthen the baseline. Likewise, scaling with system size and alternative graph topologies (rings, weak disorder, or sparse long-range couplings) will be important to assess robustness. Finally, schedule design itself is a practical lever: local adiabatic ramps, explicit end pauses, or counter-diabatic corrections could sharpen the quantum protocol without altering the problem Hamiltonian.
@@
To the best of our knowledge, this is one of the first detailed investigations of ternary optimization within a quantum annealing framework using a spin-$1$ system with a quadratic anisotropy. We find evidence for a narrow but reproducible region in the $(J,D)$ plane where quantum systems outperform classical systems under matched finite-time budgets. The advantage becomes most visible when the initial transverse scale $c$ is large and the schedule allocates sufficient time at small $g$, conditions that jointly promote coherent transport across sequences of avoided crossings near order-change boundaries, particularly in the easy-plane sector ($D>0$). Conversely, with insufficient driving or an aggressive terminal ramp, AQA can underperform even on classically easy instances.